The recursive definition of a function X is given as:
f(0)=5 and f(n)=f(n-2)+5
Now, find out the value of f(14) using the above function.
1
Expert's answer
2020-10-07T19:02:54-0400

"\mathbf{Given : f(0)=5\;and\;f(n)=f(n-2)+5}\\ \\ \\ \\ \\ 
\mathbf{To\;find:\;f(14)}"


"\mathbf{Put\;n=14\;in\;the\;recursive\;definition\;of\;function}"

"\mathbf{we \;get-}"


"\mathbf{\therefore f(14)=f(14-2)+5=f(12)+5}\\ \\
\mathbf{\implies f(14)=(f(12-2)+5)+5=f(10)+10}\\ \\
\mathbf{\implies f(14)=(f(10-2)+5)+10=f(8)+15}\\ \\
\mathbf{\implies f(14)=(f(8-2)+5)+15=f(6)+20}\\ \\
\mathbf{\implies f(14)=(f(6-2)+5)+20=f(4)+25}\\ \\
\mathbf{\implies f(14)=(f(4-2)+5)+25=f(2)+30}\\ \\
\mathbf{\implies f(14)=(f(2-2)+5)+30=f(0)+35}\\ \\
\mathbf{\implies f(14)=5+35=40\;\;(\because\; f(0)=5)}\\ \\
\mathbf{\therefore f(14)=40}"